In 36 years Ben will be 3 times as old as he is right now. How old is he right now? May you help me find the two system of equations I need?
Answer by ikleyn(52786)

.
You do not need the system of equations.

You need only ONE equation. It is enough.

The equation is

B + 36 = 3B.

The solution is B = 18.

Answer. Ben is 18 years old now.

Check. 18 + 36 = 54 = 3*18.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".